A wheel makes 71.9 rotations during 5.97s. What is the average angular velocity in that time interval?
________rad/s
Answer #1
Given that
angular displacement is θ = 71.9(2π) rad
Time taken t = 5.97s
   ------------------------------------------------------------
The average angular speed is
                                          ω = θ / t
                                            = (71.9(2π) rad ) / 5.97s
                                            = 75.67 rad/s